OpenStudy (anonymous):

Two times Owen’s age plus 2 times Charlotte’s age equals 32. Charlotte’s age is also 3 times Owen’s age. How old is Charlotte? is it 42?

OpenStudy (mokeira):

Let Owen's age be y so Charlotte's age is 3y Now with that information form an equation \[2y+2(3y)=32\] \[2y+6y=32\] \[8y=32\] \[y=4 \] so Charlotte is \[3\times 4=12years \]

OpenStudy (anonymous):

ohhh thanks!

OpenStudy (anonymous):

is this right?

OpenStudy (mokeira):

12 is th answer XD Just understand the steps

OpenStudy (anonymous):

yeah urr right thanks!